OpenStudy (goformit100):

A stone is dropped into a quiet lake and waves move in circles at the speed of 5 cm/s. At the instant when the radius of the circular wave is 8 cm, how fast is the enclosed area increasing?

OpenStudy (goformit100):

@terenzreignz

terenzreignz (terenzreignz):

These questions are eerily similar to each other :D We're looking for \[\Large \frac{dA}{dt}\] given that... \[\Large \frac{dr}{dt}=5\\\Large A= \pi r^2\] So...\[\huge \frac{dA}{dt}=\frac{dA}{dr}\times \left.\frac{dr}{dt}\right]_{r=8}\]

OpenStudy (goformit100):

Thank you sir